a metallic rod of length 2 m is rigidly clamped at its middle point.A longitudinal wave is set up in rod in such away that there will be 3 nodes on either side of clamped point.Then the wavelength of the longitudinal wave so produced is
1
Expert's answer
2012-07-17T09:45:37-0400
& the wavelength of the longitudinal wave so produced is:
l = L / n = 2 /
(3 * 2) = 0.33 m